OpenStudy (anonymous):

2x^2-17=5x write each equaton in standord form and determine (A)< (B)< and (C)

OpenStudy (anonymous):

sorry (A) (B) (C)

OpenStudy (anonymous):

there is no \[a\ , b\ , c\]

OpenStudy (anonymous):

my prob. says write each equation in standord form and determine (A), (B), and (C) 2x^2-17=5x

OpenStudy (anonymous):

quadratic equation standard form is \[ax^2+bx+c=0\]. So, what do you do to transform the expression \[2x^2-17=5x\] into \[ax^2+bx+c=0\] Check out: http://www.oojih.com/show/algebra/quadraticfunction/

OpenStudy (anonymous):

\[The \ standard\ form\ is\ 2x^2 - 5x +17 = 0\]

OpenStudy (anonymous):

Krish, not quite right

OpenStudy (anonymous):

standard form is 2x^2 - 5x -17 = 0 a=2, b= -5 and c = -17

OpenStudy (anonymous):

@chloebella Basically u hv to shift all terms to one side as i do, keeping in mind the change of sign when u shift the term after that the coefficient of x^2 term is a coefficient of x term is b and constant term is c

OpenStudy (anonymous):

Got it????? If yes, pls click GOOD ANSWER to close the question

OpenStudy (anonymous):

already gave u a medal